Clean room - Greg, Alfio

We find W#1 in clean room ready for testing.

We start attaching the network (fiber -> Mini Mc spare -> ethernet cable -> port 19 on switch 0103) and the power. Items location:
  • Power cable: metal cabinet in the electronics lab, about mid-height, in a cardboard box labeled "AGW power cord"
  • Spare Mini-MC with the fiber: AO#3 cabinet at level 3U, in a small rigid plastic bag labeled "AO WFS - AGW1&2 - MINI-MC"
  • Ethernet cable: got a long one from the electronics lab
  • Switch port: ask Stephen Hooper for configuration

W#1 functional check

We start a functional W#1 check. The cube is missing from the flowerpot, so things that need optical feedback are not tested. However all movements start up fine and no errors occur.
The network connection using the Mini-MC seems a little weak (Ethernet connector and Mini-MC power must be pressed firmly). This had already been noted in the past. If the connection goes down, all the AGW goes offline for a moment and when the network reconnects the software does not always restart correctly - known issue. As long as the network holds, everything seems to work correctly

The AGW cube is not mounted. I find a cube in the AO#3 cabinet, but it is missing the metal slit that fits into the flowerpot mount. Elliot sends a telescope-wide email to ask for informations.

NOTE: cube was found later on. See last log section.

Cube found in AO#3 cabinet. This is not the correct cube.
IMG_20160203_162259.jpg

Cooling

Elliot attaches the cooling line. The cooling water is at 12°C (so it is not the main telescope loop??) and there is no risk of condensation.

Wooden blocks

Elliot finds a few wooden blocks. We jack the AGW up and fit the blocks under it.

IMG_20160203_160447.jpg

Laser tracker setup

Greg builds a reference system in the room using four reflectors. Due to space constrains, the laser tracker head cannot be put in a position to see the whole circle on the M1 flange, but instead it must be moved in two positions to reconstruct this flange. The LUCI-facing flange instead can be measured from a single position.

The four reference points. One of them is on the high bay door (of the two doors, is the one that is opened less often), the others are on fixed places.
IMG_20160203_161935_arrows.jpg

Measurement of the two flanges.

Greg measures the reference system, then an arc of about 120° on the M1 flange. Then the tracker head is moved in the second position, the reference system measured again, and another 120° arc done.

Repeat on the LUCI side: measure the reference system, and then a full 360° arc on the flange. The orientation of both flanges is fitted by the laser tracker program and the difference is computed.

Result is 179.9983 degrees (normals are reversed), which corresponds to a difference of just 6 arcseconds. The error on these measurements is not known, Greg is working with the tracker program to calculate it.

Glycol leak

There is a smallish leak (a few drops per minute) from where the cooling lines connect to the AGW. We are at the end of the afternoon and we can close the lines (the AGW can stay off even if we want to go on with the laser tracker). Elliot is going to fix it later today.

Leak temporary fix: IMG_20160203_174758.jpg

Cube found

Michael Lefebvre informs us that the cube is in a cabinet inside the high bay container. There we find both cubes (DX and SX). After the cube is mounted, I follow the "W_unit_functional_test_vB.docx" in the handover documents sections. The document was written for the DX system and some small changes are noted. All devices pass the functional test, including camera lens and the three tip-tilt axes.

CCD39 noise is very good:
------------------------------ SUMMARY ------------------------------------
Bin: 1 - Framerate  990 Hz (2500 kpix/sec) - RON (e-):   8.61 (estimated)
Bin: 2 - Framerate  990 Hz ( 800 kpix/sec) - RON (e-):   6.37 (estimated)
Bin: 2 - Framerate  625 Hz ( 380 kpix/sec) - RON (e-):   4.00 (estimated)
Bin: 2 - Framerate  400 Hz ( 380 kpix/sec) - RON (e-):   3.98 (estimated)
Bin: 3 - Framerate  200 Hz ( 380 kpix/sec) - RON (e-):   4.08 (estimated)
Bin: 4 - Framerate  300 Hz ( 335 kpix/sec) - RON (e-):   4.19 (estimated)
Bin: 4 - Framerate  200 Hz ( 335 kpix/sec) - RON (e-):   4.25 (estimated)
Bin: 4 - Framerate  105 Hz ( 335 kpix/sec) - RON (e-):   4.31 (estimated)

Pupil wobbling measurement gives a somewhat high result (5 pix), but only one pupil was correctly centered and the result might not be precise. This measure must be repeated with a better centering.

20160203_pupilwobble.png

-- AlfioPuglisi - 03 Feb 2016
Topic revision: r2 - 04 Feb 2016, AlfioPuglisi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback